OpenStudy (geometry_hater):

For f(x) 3x - x^2 Find: f(x+h) f(x) ---------- h

OpenStudy (anonymous):

Is there a picture? (I don't understand the question.)

OpenStudy (anonymous):

i think it should be: [f(x+h)-f(x)]/h

OpenStudy (anonymous):

then f(x+h)=3(x+h)-(x+h)^2 and f(x)=3x-x^2 and you make the substitution, crank through the algebra, cancel the h's in the appropriate places, and you are done

OpenStudy (geometry_hater):

Oh ok i was just confused over the h part. The substitution i got hands down.

OpenStudy (anonymous):

okeedokee.